OpenStudy (pphalke):

two planes leave an airport at the same time, one flying east, the other flying west. the eastbound plane travels 100 mph slower. they are 4000 mi apart after 5 hr. find the speed of each plane. @janu16

OpenStudy (janu16):

5x + 5(x+100) = 4000 x=350 I defined x as the speed of the slower speed so the slower plane has speed of 350 mph I defined x+100 as the speed of the faster plane, so faster plane goes (350+100) or 450 mph check: what is 4(330) + 4(430)=4000

OpenStudy (janu16):

slower plane is 350mph faster plane is 450 mph

OpenStudy (janu16):

@pphalke sorry i took long

OpenStudy (janu16):

@pphalke do you understand that?

OpenStudy (anonymous):

rate times time equals distance. { 5 west + 5 east = 4000, east = west - 100 } // Solve \[{ east = 350, west = 450 }\]